OpenStudy (anonymous):

medal will be given find f(-2)

OpenStudy (anonymous):

f(-2) means that you are meant to evaluate a function at the f(x), where f is any function. We cannot find the f(-2) for you if we don't know the function. For example If f(x)=2x+3 Then for the f(-2), we would simply replace all instances of x with -2 f(-2)=2(-2)+3=-1 If f(x)=x*x+6, then f(-2)=(-2)*(-2)+6=10